# Optimize the Unseen - Fast NeRF Cleanup with Free Space Prior

### v1

@@ 7.1 Method Robustness
In Figure 9, and the attached videos, we compare our method to the Nerfacto baseline. The results confirm that our method effectively removes artifacts while preserving scene details across a wide range of scenes, underscoring its robustness beyond the Nerfbusters dataset.
7.2 Full Quantitative Results
Table 1 presents the complete evaluation results for all methods, directly correlating with the points depicted in Figure 4 of the main paper. Each row in Table 1 corresponds to a point in the PSNR vs. Coverage plot. Beyond the standard PSNR and Coverage metrics shown in the main paper, we also report SSIM, LPIPS, and Dice scores for a more comprehensive evaluation. While SSIM and LPIPS have been previously evaluated for BayesRays, we introduce the Dice score to provide additional insights into the cleanup results.
@@ 7.2 Full Quantitative Results
The Dice score offers a nuanced perspective that the coverage metric alone cannot provide. Coverage measures how completely a method reconstructs regions visible from the training cameras but does not penalize excessive retention of artifacts. In contrast, the Dice score penalizes methods that inaccurately retain content in unseen areas. For example, as shown in Table 1, the baseline achieves a high coverage score because it does not perform any cleanup. However, Table 1 also reveals a low Dice score for the baseline, highlighting its inability to remove artifacts from unseen regions effectively.
@@
7.3 Additional Qualitative Results
We present qualitative results on additional scenes from the Nerfbusters dataset in Figure 10. These examples further support our claim of achieving a balance between effective cleanup and scene preservation. Our method demonstrates its ability to remove artifacts while maintaining the integrity of the reconstructed regions, even in challenging scenarios.
Furthermore, the qualitative results highlight the limitations of Gaussian Splatting. Notably, it exhibits significant noise in novel views located far from the training cameras. This issue aligns with the quantitative findings, as the excessive noise contributes to the low PSNR scores reported in Table 1. These observations underline the importance of robust artifact removal in achieving high-quality reconstructions for novel viewpoints.

### v3

@@ A.3 Method Robustness
In Fig.˜9, and the attached videos, we compare our method to the Nerfacto baseline. The results confirm that our method effectively removes artifacts while preserving scene details across a wide range of scenes, underscoring its robustness beyond the Nerfbusters dataset.
A.4 Evaluation Using the Old Paradigm
As stated in the main paper, the conventional metric comparison has relied on GT masks to measure PSNR and Coverage. This approach allows methods that do not clean regions outside the area of interest to achieve high PSNR and, more importantly, artificially inflated Coverage. In Fig.˜10, we demonstrate that this old paradigm introduces a significant bias in favor of the baseline, despite its practical noisiness, as it disregards artifact-heavy regions outside the GT masks. Moreover, we show that our method achieves SOTA performance even under the conventional paradigm, while also excelling in our upgraded metric paradigm.
A.5 Choosing $\lambda$
Our method is robust to a wide range of $\lambda$ values as can be seen in 4, $\lambda\in[10^{-5},1]$ yields both high PSNR (17.35–17.69) and high coverage (87.57%–96.45%). This robustness stems from our method’s design, which optimizes regions where the reconstruction loss is inactive. Even with small $\lambda$ values, the FSP loss dominates in these regions, yielding much cleaner results than default Nerfacto. We did not include experiments with $\lambda>1$, as this would imbalance the losses and violate the principle that the FSP loss should complement, not compete with, the reconstruction loss.
A.6 Full Quantitative Results
Tab.˜4 presents the complete evaluation results for all methods, directly correlating with the points depicted in Figure 4 of the main paper. Each row in Tab.˜4 corresponds to a point in the PSNR vs. Coverage plot. Beyond the standard PSNR and Coverage metrics shown in the main paper, we also report SSIM, LPIPS, and Dice scores for a more comprehensive evaluation. While SSIM and LPIPS have been previously evaluated for BayesRays, we introduce the Dice score to provide additional insights into the cleanup results.
Figure 10: Evaluation using old paradigm: Comparison of our method, BayesRays, and the baseline on the Nerfbusters dataset, calculating PSNR solely within the GT masks.
@@ A.6 Full Quantitative Results
Figure 12: Density Histogram Underwater: Histogram of densities across the Curacao scene after sigmoid-softened, shifted to [0, 1]. (Left) Histogram of densities across the scene before the cleanup process. (Right) Histogram of the densities across the scene after the cleanup process.
The Dice score offers a nuanced perspective that the coverage metric alone cannot provide. Coverage measures how completely a method reconstructs regions visible from the training cameras but does not penalize excessive retention of artifacts. In contrast, the Dice score penalizes methods that inaccurately retain content in unseen areas. For example, as shown in Tab.˜4, the baseline achieves a high coverage score because it does not perform any cleanup. However, Tab.˜4 also reveals a low Dice score for the baseline, highlighting its inability to remove artifacts from unseen regions effectively.
@@
A.7 Additional Qualitative Results
We present qualitative results on additional scenes from the Nerfbusters dataset in Fig.˜11. These examples support our claim of achieving a balance between effective cleanup and scene preservation.
Furthermore, the qualitative results highlight the limitations of Gaussian Splatting. Notably, it exhibits significant noise in novel views located far from the training cameras. This issue aligns with the quantitative findings, as the excessive noise contributes to the low PSNR scores reported in Tab.˜4. These observations underline the importance of robust artifact removal in achieving high-quality reconstructions for novel viewpoints.
@@ A.7 Additional Qualitative Results
Figure 13: Noisy input images study: (Top) Noised input image. (Middle) Original images and the novel view from the Nerfacto baseline. (Bottom) Original images and the novel view after cleanup.
A.8 Noisy Images Study
As stated in our limitations section, our approach is based on the reconstruction loss, which relays on a good data. To demonstrate this, we added gaussian noise to the input images of a scene, build a NeRF and applied our method on the noisy scene. As can be seen in Fig.˜13, the cleanup did work on under-optimized regions, but did not fix the color bias caused by the noisy input images.
A.9 Densities in Complex Medium
As seen in Fig.˜1, our approach works on underwater scenes like Seathru dataset. But the assumption that "There is nothing in the unseen area" does not mean densities should be zero. This is because the physics of underwater imaging involve attenuation and back scattering, which means density should not be zero.
We investigate this further by calculating a density histogram for the underwater scene Curacao from the Seathru dataset. See Fig.˜12. As can be seen, the original density histogram (left) is pushed towards the bi-modal histogram (right) that we find in regular images (cf. see Fig.˜6). Because the prior does not match the imaging physics the resulting histogram is not bi-modal, leaving enough non-zero densities to handle the image properly, as shown visually by the results on the Seathru dataset.